OpenStudy (idealist10):

Solve (sec^2 y)y'-3tan(y)=-1.

OpenStudy (xapproachesinfinity):

rewrite it as tany first

OpenStudy (idealist10):

What do you mean? How?

OpenStudy (xapproachesinfinity):

oh hold on! add 3tany to both sides and then divide by sec^2y

OpenStudy (anonymous):

Substitute \(u=\tan y\), then \(u'=\sec^2y~y'\). The DE then changes to \[u'-3u=-1\]

OpenStudy (xapproachesinfinity):

it is a lot easier the way siths did it!

OpenStudy (anonymous):

Or treat it as a separable equation: \[\begin{align*} \sec^2y~y'&=3\tan y-1\\ \frac{\sec^2y}{3\tan y-1}~dy&=dx \end{align*}\] Then integrate with a substitution, \(u=3\tan y-1\), so \(\dfrac{1}{3}du=\sec^2y~dy\): \[\frac{du}{3u}=dx\]

OpenStudy (idealist10):

But what's the next step of u'-3u=-1?

OpenStudy (xapproachesinfinity):

that's what i was thinking about siths^_^

OpenStudy (xapproachesinfinity):

have you ever done DE?

OpenStudy (idealist10):

Yes.

OpenStudy (anonymous):

Well my first suggested method changes the equation to a linear equation, so you can find the integrating factor and so on. The second method is simpler (I guess you could call it that) because separable equations are generally easy to solve.

OpenStudy (idealist10):

So du/3u=dx after integrating, you would get ln abs(3u)=x, right?

OpenStudy (anonymous):

+C, yes

OpenStudy (idealist10):

Got it! Thanks!

OpenStudy (anonymous):

yw